Locate the absolute maximum and minimum for each of the following functions and justify your responses.

Question 1: 𝑓(π‘₯) = cube root of x^2 on the interval [βˆ’1,1]Β 

Question2: 𝑓(π‘₯)= 𝑔(π‘₯) = π‘₯𝑒^2π‘₯ on the interval [βˆ’2,0]Β 


Question 3: Given 𝑓(π‘₯) = 3π‘₯^3 βˆ’ 18π‘₯^2 βˆ’ 45π‘₯ + 10

a) On what interval(s), is 𝑓(π‘₯) increasing? Justify.

b) what value(s) of π‘₯ does 𝑓(π‘₯) have a relative minimum?

c) On what interval(s), is 𝑓(π‘₯) decreasing and concave up? Justify.


1
Expert's answer
2021-02-03T01:15:32-0500

Question 1:

"f(x)=(x^2)^{\frac{1}{3}}" on the interval [βˆ’1,1]


"f'(x)=\frac{2}{3}x^{-\frac{1}{3}}"




"f'(x)<0" when "x<0" and "f'(x)>0" when "x>0"

"f'(x)" does not exists when "x=0"

"f(-1)=f(1)=1" - absolute maximum on the interval [βˆ’1,1]

"f(0)=0" - absolute minimum on the interval [βˆ’1,1]


Question2: "g(x)=xe^{2x}" on the interval [βˆ’2,0]


"g'(x)=e^{2x}+2xe^{2x}=(1+2x)e^{2x}"

"g'(x)=0" at the point "x=-\frac{1}{2}" . This is point of minimum because "g'(x)<0" when "x<-\frac{1}{2}" ,

and "g'(x)>0" when "x>-\frac{1}{2}" .

"g(-\frac{1}{2})=-\frac{1}{2}e^{-1}" - absolute minimum on the interval [βˆ’2,0].


"g(-2)=-2e^{-4}" and "g(0)=0"

"g(0)>g(-2)" , so g(0)=0 - absolute maximum on the interval [βˆ’2,0].




Question 3: Given "f(x)=3x^3-18x^2-45x+10"

a) On what interval(s), is 𝑓(π‘₯) increasing? Justify.


"f'(x)=9x^2-36x-45=9(x^2-4x-5)=9(x+1)(x-5)"

"f'(x)>0" on the intervals "(-\infty, -1)" and "(5,+\infty)" , so "f(x)" is increasing on these intervals


b) what value(s) of π‘₯ does 𝑓(π‘₯) have a relative minimum?


at the point "x=5" function "f(x)" has a relative minimum because "f'(x)" changes its sign from

minus to plus


c) On what interval(s), is 𝑓(π‘₯) decreasing and concave up? Justify.


"f'(x)<0" on the interval "(-1,5)" , so "f(x)" is decreasing on this interval.


"f''(x)=18x-36=18(x-2)"

"f''(x)>0" on the interval "(2, +\infty)" , so the function is concave up on this interval